m doing Variables and i just cant get this problem c=2(pie)R
Found 2 solutions by checkley77, rfer:
Answer by checkley77(12569)

c=2(pie)R
This formula is for finding the circumferance (c) of a circle.
Pie=3.14
R=the radius or diameter/2 of the circle.
Example:
R=7
c=2*3.14*7=43.96 is the circumferance.

Answer by rfer(10417)

C is the circumference of a circle
r is the radius
d is diameter
pi =3.14 or 22/7
the formula for circumference is C=d*pi, or C=2r*pi
Does that help?
Bob